OpenStudy (anonymous):

A line has a y-intercept of 4 and a slope of 8/7. Find the equation of the line. A. y = 4x + 8/7 B. y = 4x - 7/8 C. y = 8/7x + 4 D. y = 7/8x - 4

OpenStudy (lucaz):

y-intercept of 4 means when x=0 y=4, you have a point (0,4), and the slope 8/7

OpenStudy (lucaz):

use the point-slope form \[\Large{ y-y_1=m(x-x_1) }\]

OpenStudy (lucaz):

\[\large{ y-(4)=\frac{8}{7}(x-0)}\]

OpenStudy (anonymous):

thanks :D

OpenStudy (lucaz):

no problem